OpenStudy (anonymous):

What is the slope of the line parallel to the line 3x-2y=12 ? A. -2 B. -2/3 C. 3/2 D. 1/4

OpenStudy (the_fizicx99):

In y = mx + b The slope is M The y intercept is b 3x-2y=12 Subtract both sides by 3x -2y = -3x + 12 Divide both sides by -2 to isolate y y = 3/2x - 6

OpenStudy (anonymous):

So the parallel line would have the same slop right? 3/2 ?

OpenStudy (the_fizicx99):

Yea

OpenStudy (anonymous):

Thank you!!!!! (:

OpenStudy (the_fizicx99):

The slope is M :3